calculate the molarity of the NaCl solution when 29.2 g of NaCl is dissolved in 100 mL of water
1
Expert's answer
2015-11-24T14:46:10-0500
If we assume that dissolution does not affect the volume of solution, C = n/V,
n = m/MW
C = m/(MW V) = 29.2 g/(58.44 g mol-1 100 mL) = 4.996 M ≈ 5.00 M

Answer: 5.00 M